# Graph Neural Network-Based Reinforcement Learning for Controlling Biological Networks – the GATTACA Framework

### v2

@@ VI-B Performance Evaluation Methodology
We compare the results of the GATTACA framework with those obtained using the state-of-the-art exact attractor-based sequential instantaneous source-target control (ASI) algorithm introduced in [39] and implemented in the CABEAN software tool [5]. The exact algorithm follows the divide-and-conquer approach based on SCC-decomposition of BN structure graph and it leverages symbolic techniques based on Binary Decision Diagrams (BDDs) to encode the BN dynamics. Amongst the control algorithms provided by CABEAN, we selected ASI as the best match for the control problem considered in this study.
@@
However, in the case of all environmental conditions considered, e.g. the T-LGL model, fully specified environmental conditions are enumerated and considered one by one in CABEAN. For each such condition, all attractors are computed with CABEAN and divided into source attractors, i.e. those not aligned with the target configuration, and target attractors, i.e. those aligned with the target configuration. Then, for each source attractor, the shortest path to any of the target attractors is computed with CABEAN. If there is no available target attractor for a given environmental condition, we check all other environmental conditions for a target attractor. We first switch the environmental condition, and then (if necessary) use CABEAN to control the network within the new environmental condition (the GATTACA framework can switch environmental conditions automatically by letting it perturb input nodes). We provide an example of this scenario in Appendix -B. To account for the “manual” environmental condition switch, one is added to the control path length returned by CABEAN. We take the minimum length over all other environmental conditions containing a target attractor. Finally, the averaged length of the shortest paths over all environmental conditions and source attractors is reported.
Formally, the average control path length determined by CABEAN for a BN model $M$, denoted $ACPL_{CABEAN}(M)$, is defined as follows:
{strip}
$\text{ACPL}_{\mathrm{CABEAN}}(M)\;=\;\mathbb{E}_{e\in E_{M}}\Biggl{[}\,\mathbb{E}_{s\in\mathcal{S}^{M}_{e}}\Biggl{[}\begin{cases}\displaystyle\min_{t\in\mathcal{T}^{M}_{e}}\!\text{cpl}(s,t)&\text{if }\mathcal{T}^{M}_{e}\neq\emptyset,\\[4.30554pt] \displaystyle\min_{e^{\prime}\in E_{M}/\{e\}}\Bigl{(}1+\min_{t\in\mathcal{T}^{M}_{e^{\prime}}}\!\text{cpl}(s,t)\Bigr{)}&\text{otherwise.}\end{cases}\Biggr{]}\Biggr{]},$ (2)
where $E_{M}$ is the set of all separate environmental conditions considered for model $M$, $\mathcal{S}^{M}_{e}$ is the set of source attractors, $\mathcal{T}^{M}_{e}$ is the set of target attractors for a given environmental condition $e$ for model $M$, and cpl is the control path length computed by CABEAN for a specific pair of source and target attractors. Notice that in the case of a fully or partially specified environmental condition where all the corresponding fully specified conditions can be handled by CABEAN at once, the set $E_{M}$ in Equation (2) is a singleton. In the case where all environmental conditions are considered, the cardinality of $E_{M}$ is equal to the number of distinct possible settings of the values for all input nodes.
@@ VII-A Evaluation of iPASIP
Our iPASIP method, although it is a purely statistical method based on model simulations, provides a strong alternative to traditional methods for finding BN attractor states. Table III presents summarised results for each model where CABEAN was able to compute the attractors for the environmental conditions specified in Table II.
@@
In the case of the MAPK model, CABEAN failed in 3 of 16 possible environmental conditions. Running iPASIP under all environmental conditions resulted in a very large number of identified PA states, which we could not verify due to the lack of ground-truth information. However, running iPASIP under all environmental conditions except the ones in which CABEAN failed, resulted in fully correct identification of all attractors, see Table III.
@@ F-A C-A Spurious F-A C-A Failed
For the Bladder model, iPASIP found 25 PA states in total in all the 16 environmental conditions, while CABEAN returned 24 attractors in 15 environmental conditions and failed in one, namely where the EGFR_stimulus input node is set to 1 and all the remaining three input nodes are set to 0. All the 24 CABEAN-found attractors are represented within the 25 PA states. By computing an attractor reachable from the extra PA state identified by iPASIP, we found that it is in fact an attractor state of an attractor consisting of 184,320 states under the environmental condition that CABEAN was unable to handle.
@@
In summary, the iPASIP approach provides a scalable and effective solution for identifying true PA states, as defined in Definition III.1, in large-size BN models. Moreover, as the case studies show, iPASIP often successfully identifies all attractor states of a large BN model.
@@ VII-B Evaluation of the GATTACA framework
With regard to control problem solutions found by our GATTACA framework, the average length of the control strategies is typically within one control action/intervention of the optimal solution computed in accordance with the formula in Equation (2) based on the control path lengths obtained by the exact ASI algorithm implemented in the CABEAN software tool, as shown in Table IV.
However, for the ABA model the number of environmental conditions is too big ($2^{23}$) to be handled by CABEAN (entries marked as N/A in Table IV). As obtained with the GATTACA framework, the ABA model is characterised by a very large number of PA states. Nevertheless, the GATTACA framework successfully identifies effective control strategies, demonstrating the scalability of our framework in this respect. Moreover, we conducted experiments in which the number of PA states was overestimated by a factor of 1000 and ATGC was still able to identify effective control strategies (data not shown). Those two facts prove that the ATGC is resilient to both high attractor counts and spurious PA states.
@@
To further investigate the scalability potential of the GATTACA framework, we analysed the MCF-7 model under all possible environmental conditions. As in the case of the ABA model, MCF-7 is characterised by a large number of PA states and the GATTACA framework is still capable of successfully identifying control strategies as shown in Table V. Unfortunately, we were unable to obtain the exact results where all $256$ fully specified environmental conditions were considered with CABEAN as it could not finish the computations of attractors within 168 hours (one week) on an 8-Core 2.25 GHz AMD EPYC 7742 processor with 256 GB of RAM.
Next, we analysed the original T-diff model with 71 nodes, introduced in [36], under two environmental conditions: 1) the original, partially specified (PS) condition from [36] and 2) the fully specified (FS) condition from [25] presented in Table II. The former condition leaves some of the input nodes unspecified with respect to the latter, i.e. IFNGR1, IFNGR2, GP130, IL6RA, CGC, IL12RB2, IL10RB, IL10RA, IL15RA, IL2RB, and IL27RA.
The original 71-node T-diff model, under both conditions, is characterised by a large number of PA states identified by iPASIP, i.e. 156,864 and 25,856 states, respectively. CABEAN fails to identify the attractors and cannot compute control strategies for these BNs. However, the GATTACA framework successfully identifies control strategies in both cases. The average control strategy length is 6.17, as shown in Table V. It is noticeably larger than for all other BN models considered in this study. Nevertheless, this is justifiable, as such large numbers of PA states suggest that the two BN models have large numbers of attractors. Meanwhile, the size of the original T-diff model (71 nodes) is comparable to the size of the model in [25] (68 nodes), the latter with only 12 attractors; see Table IV. It is then highly probable that the state space of the original T-diff model under the two environmental conditions is divided into much finer-grained basins of attraction than in the case of the modified model in [25]. Consequently, driving the dynamics towards a PA state aligned with the given target configuration requires an action that places the network dynamics into a relatively small basin of attraction, likely necessitating simultaneous perturbation of many genes. Since this is restricted due to the limit imposed on the maximum number of genes that can be perturbed at once within a single ATGC action, identified control strategies must traverse multiple attractors to reach the target configuration.
@@
TABLE V: Average control strategy lengths obtained with the GATTACA framework for the MCF-7 model under all environmental conditions and the original 71-node T-diff model of [36] under two environmental conditions (env. cond.), i.e. the partially specified (PS) one in [36] and the fully specified (FS) one in [25]. Both conditions are characterised by large number of PA states identified by iPASIP.
Finally, we performed one additional scalability test of the GATTACA framework on a randomly generated BN model of 200 nodes (Random200). The model is challenging as it has no input nodes and its structure graph consists of a single SCC. Therefore, the attractor detection algorithm and the ASI control algorithm in CABEAN cannot exploit the SCC-decomposition and CABEAN fails to compute the control strategies within 168 hours (one week). The GATTACA framework identifies 165 PA states and returns control strategies of average length of 7.1 for the target configuration $x_{3}=1$, where $x_{3}$ is one of the nodes; see Table IV. This relatively long control strategy on average can be explained by the fact that a network of highly interconnected nodes may require a simultaneous perturbation of a larger number of nodes than in the case of a loosely connected network. Similarly to the two cases of the original T-diff model in Table V, since we limit the number of simultaneous perturbations, the GATTACA framework finds longer control strategies that traverse a number of intermediate PA states.
@@ VIII Conclusions
We formulate the attractor-target control problem in the context of cellular reprogramming – a novel control problem for biological networks modelled using the BN framework under the asynchronous update mode. To provide a scalable solution, we develop and implement the GATTACA framework. Our framework can handle models of sizes that are prohibitive for the existing state-of-the-art control algorithms.
@@
Moreover, the scalability is achieved by considering previously introduced concept of a pseudo-attractor and by exploiting iPASIP, which we develop by improving previously proposed approach for effective identification of PA states in large BN models. We demonstrate the effectiveness of iPASIP on a number of BN models of real-world biological networks of different sizes by comparing the identified PA states with true attractors obtained with an exact state-of-the-art algorithm implemented in the CABEAN software tool. We show that iPASIP is capable of correctly identifying representative attractor states of all attractors in most of the cases. It also manages to detect attractor states in BN models for which CABEAN fails to compute the attractors. Furthermore, iPASIP proves its scalability and effectiveness as part of the GATTACA framework, where the determined PA states facilitate the computation of control strategies for large BN models of various real-world biological networks.
@@
The proposed GATTACA framework offers a scalable and effective solution to the attractor-target control problem as demonstrated by he results of the conducted computational experiments on a number of large real-world biological networks from literature. Given source attractor state and target configuration of genes, the ATGC of the GATTACA framework finds proper control strategies that drive the network from the source pseudo-attractor to the target configuration by taking actions (intervening) only in intermediate PA states that correspond to phenotypical cellular states that can be observed in the lab. Our framework is capable of finding optimal or suboptimal control strategies which lengths in most of the cases differ on average from the optimal strategies by at most one intervention. Moreover, the GATTACA framework is capable of identifying effective control strategies of reasonable lengths in BN models on which CABEAN fails. These models are large in terms of number of nodes and often characterised by many (pseudo-)attractor states. An additional evaluation conducted on a challenging network which structure graph is a single large SCC of 200 nodes, which analysis is beyond the capabilities of state-of-the-art exact solutions, further demonstrates the unprecedented scalability potential of the GATTACA framework in controlling large models of complex biological networks.
IX Acknowledgments
@@
This research was funded in whole or in part by the National Science Centre, Poland under the OPUS call in the Weave programme, grant number 2023/51/I/ST6/02864. For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.

### v3

@@ VI-B Performance Evaluation Methodology
We compare the results of the GATTACA framework with those obtained using the state-of-the-art exact attractor-based sequential instantaneous source-target control (ASI) algorithm introduced in [cmsb19] and implemented in the CABEAN software tool [SP21]. The exact algorithm follows the divide-and-conquer approach based on SCC-decomposition of BN structure graph and it leverages symbolic techniques based on Binary Decision Diagrams (BDDs) to encode the BN dynamics. Amongst the control algorithms provided by CABEAN, we selected ASI as the best match for the control problem considered in this study.
@@
However, in the case of all environmental conditions considered, e.g. the T-LGL model, fully specified environmental conditions are enumerated and considered one by one in CABEAN. For each such condition, all attractors are computed with CABEAN and divided into source attractors, i.e. those not aligned with the target configuration, and target attractors, i.e. those aligned with the target configuration. Then, for each source attractor, the shortest path to any of the target attractors is computed with CABEAN. If there is no available target attractor for a given environmental condition, we check all other environmental conditions for a target attractor. We first switch the environmental condition, and then (if necessary) use CABEAN to control the network within the new environmental condition (the GATTACA framework can switch environmental conditions automatically by letting it perturb input nodes). We provide an example of this scenario in Appendix -B. To account for the “manual” environmental condition switch, one is added to the control path length returned by CABEAN. We take the minimum length over all other environmental conditions that contain a target attractor. Finally, the averaged length of the shortest paths is reported over all environmental conditions and source attractors.
Formally, the average control path length for a BN model $M$, calculated from the exact results obtained with CABEAN, is denoted $\mathsf{ACPL}_{\mathrm{CABEAN}}(M)$ and is defined in accordance with the following formula:
where $E_{M}$ is the set of all separate environmental conditions considered for model $M$, $\mathcal{S}^{M}_{e}$ is the set of source attractors, $\mathcal{T}^{M}_{e}$ is the set of target attractors for a given environmental condition $e$ for model $M$, and $\mathsf{cpl}(s,t)$ is the control path length computed by CABEAN for a specific pair of source ($s$) and target ($t$) attractors. Notice that in the case of a fully or partially specified environmental condition where all the corresponding fully specified conditions can be handled by CABEAN at once, the set $E_{M}$ in Equation (LABEL:eq:acpl) is a singleton. In the case where all environmental conditions are considered, the cardinality of $E_{M}$ is equal to the number of distinct possible settings of the values for all input nodes.
The remaining parameters of the GATTACA framework are set as follows: Adam lr = $1\mathrm{e}{-4}$, $\gamma=0.99$, batch size = 128, replay = $10^{6}$, PER $\alpha$ = 0.6, $\beta=0.4\to 1.0$, target update $\tau=0.01$, $\epsilon$-greedy ($1.0\to 0.05$, $10^{6}$ steps), and gradient clip = 10.
@@ VII-A Evaluation of iPASIP
Our iPASIP method, although it is a purely statistical method based on model simulations, provides a strong alternative to traditional methods for finding BN attractor states. Table III presents the summarised results for each model in which CABEAN was able to compute the attractors for the environmental conditions specified in Table II.
@@
In the case of the MAPK model, CABEAN failed in 3 of 16 possible environmental conditions. Running iPASIP under all environmental conditions resulted in a very large number of identified PA states, which we could not verify due to the lack of ground-truth information. However, running iPASIP under all environmental conditions except the ones in which CABEAN failed resulted in fully correct identification of all attractors, see Table III.
@@ F-A C-A Spurious F-A C-A Failed
For the Bladder model, iPASIP found 25 PA states in total in all 16 environmental conditions, while CABEAN returned 24 attractors in 15 environmental conditions and failed in one, namely where the EGFR_stimulus input node is set to 1 and all the remaining three input nodes are set to 0. The 24 CABEAN-found attractors are represented within the 25 PA states. By computing an attractor reachable from the extra PA state identified by iPASIP, we found that it is in fact an attractor state of an attractor consisting of 184,320 states under the environmental condition that CABEAN was unable to handle.
@@
In summary, the iPASIP approach provides a scalable and effective solution for identifying true PA states, as defined in Definition III.1, in large-size BN models. Moreover, as the case studies show, iPASIP often successfully identifies all the attractor states of a large BN model.
@@ VII-B Evaluation of the GATTACA framework
With regard to control problem solutions found by our GATTACA framework, the average length of the control strategies is typically within one control action/intervention of the optimal solution computed in accordance with the formula in Equation (LABEL:eq:acpl) based on the control path lengths obtained by the exact ASI algorithm implemented in the CABEAN software tool, as shown in Table IV.
However, for the ABA model the number of environmental conditions is too large ($2^{23}$) to be handled by CABEAN (entries marked N/A in Table IV). As obtained with the GATTACA framework, the ABA model is characterised by a very large number of PA states. However, the GATTACA framework successfully identifies effective control strategies, demonstrating the scalability of our framework in this respect. Moreover, we conducted experiments in which the number of PA states was overestimated by a factor of 1000 and ATGC was still able to identify effective control strategies (data not shown). Those two facts prove that the ATGC is resilient to both high attractor counts and spurious PA states.
@@
To further investigate the scalability potential of the GATTACA framework, we analysed the MCF-7 model under all possible environmental conditions. As in the case of the ABA model, MCF-7 is characterised by a large number of PA states and the GATTACA framework is still capable of successfully identifying control strategies, as shown in Table V. Unfortunately, we were unable to obtain exact results where all $256$ fully specified environmental conditions were considered with CABEAN as it could not complete the attractors calculations in 168 hours (one week) on an 8-Core 2.25 GHz AMD EPYC 7742 processor with 256 GB of RAM.
Next, we analysed the original T-diff model with 71 nodes, introduced in [tdiff], under two environmental conditions: 1) the original, partially specified (PS) condition from [tdiff] and 2) the fully specified (FS) condition from [SP20a] presented in Table II. The former condition leaves some of the input nodes unspecified with respect to the latter, i.e. IFNGR1, IFNGR2, GP130, IL6RA, CGC, IL12RB2, IL10RB, IL10RA, IL15RA, IL2RB, and IL27RA.
The original 71-node T-diff model, under both conditions, is characterised by a large number of PA states identified by iPASIP, i.e. 156,864 and 25,856 states, respectively. CABEAN fails to identify the attractors and cannot compute control strategies for these BNs. However, the GATTACA framework successfully identifies control strategies in both cases. The average control strategy length is 6.17, as shown in Table V. It is noticeably larger than for all other BN models considered in this study. Nevertheless, this is justifiable, as such large numbers of PA states suggest that the two BN models have large numbers of attractors. Meanwhile, the size of the original T-diff model (71 nodes) is comparable to the size of the model in [SP20a] (68 nodes), the latter with only 12 attractors; see Table IV. The state space of the original T-diff model under the two environmental conditions is then highly probable to be divided into much finer-grained basins of attraction than in the case of the modified model in [SP20a]. Consequently, driving the dynamics towards a PA state aligned with the given target configuration requires an action that places the network dynamics into a relatively small basin of attraction, likely necessitating simultaneous perturbation of many genes. Since this is restricted due to the limit imposed on the maximum number of genes that can be perturbed at once within a single action, identified control strategies must traverse multiple attractors to reach the target configuration.
@@
TABLE V: Average control strategy lengths obtained with the GATTACA framework for the MCF-7 model under all environmental conditions and the original 71-node T-diff model of [tdiff] under two environmental conditions (env. cond.), i.e. the partially specified (PS) one in [tdiff] and the fully specified (FS) one in [SP20a]. Both conditions are characterised by large number of PA states identified by iPASIP.
Finally, we performed one additional scalability test of the GATTACA framework on a randomly generated BN model of 200 nodes (Random200). The model is challenging as it has no input nodes and its structure graph consists of a single SCC. Therefore, the attractor detection algorithm and the ASI control algorithm in CABEAN cannot exploit the SCC-decomposition and CABEAN fails to compute the control strategies within 168 hours (one week). The GATTACA framework identifies 165 PA states and returns control strategies of average length of 7.1 for the target configuration $x_{3}=1$, where $x_{3}$ is one of the nodes; see Table IV. This relatively long control strategy on average can be explained by the fact that a network of highly interconnected nodes may require a simultaneous perturbation of a larger number of nodes than in the case of a loosely connected network. Similarly to the two cases of the original T-diff model in Table V, since we limit the number of simultaneous perturbations, the GATTACA framework finds longer control strategies that traverse a number of intermediate PA states. Training of the GATTACA framework on the 200-node network required approximately 2.5h (A100) and 0.6M steps.
@@ VIII Conclusions
We formulate the attractor-target control problem in the context of cellular reprogramming – a novel control problem for biological networks modelled using the BN framework under the asynchronous update mode. To provide a scalable solution, we develop and implement the GATTACA framework. Our framework can handle models of sizes that are prohibitive for existing state-of-the-art control algorithms.
@@
Moreover, scalability is achieved by considering the previously introduced concept of a pseudo-attractor and by exploiting iPASIP, which we develop by improving previously proposed approach for effective identification of PA states in large BN models. We demonstrate the effectiveness of iPASIP on a number of BN models of real-world biological networks of different sizes by comparing the identified PA states with true attractors obtained with an exact state-of-the-art algorithm implemented in the CABEAN software tool. We show that iPASIP is capable of correctly identifying the representative attractor states of all attractors in most of the cases. It also manages to detect attractor states in BN models for which CABEAN fails to compute the attractors. Furthermore, iPASIP proves its scalability and effectiveness as part of the GATTACA framework, where the determined PA states facilitate the computation of control strategies for large BN models of various real-world biological networks.
@@
The proposed GATTACA framework offers a scalable and effective solution to the attractor-target control problem, as demonstrated by the results of the conducted computational experiments on a number of large real-world biological networks from the literature. Given the source PA state and target configuration of genes, the ATGC of the GATTACA framework finds proper control strategies that drive the network from the source PA to the target configuration by taking actions (intervening) only in intermediate PA states that correspond to phenotypical cellular states that can be observed in the lab. Our framework is capable of finding optimal or close to optimal control strategies, where in the latter case the lengths in most of the cases differ on average from the optimal strategies by at most one intervention. Moreover, the GATTACA framework is capable of identifying effective control strategies of reasonable lengths in BN models on which CABEAN fails. These models are large in terms of number of nodes and often characterised by many (pseudo-)attractor states. An additional evaluation conducted on a challenging network which structure graph is a single large SCC of 200 nodes, the analysis of which is beyond the capabilities of state-of-the-art exact solutions, further demonstrates the unprecedented scalability potential of the GATTACA framework in controlling large Boolean models of complex biological networks.
Acknowledgments
@@
This research was funded in whole or in part by the National Science Centre, Poland under the OPUS call in the Weave programme, grant number 2023/51/I/ST6/02864.
For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
-A Example: Boolean network
Let us consider a Boolean Network $\textsf{BN}=(\mathbf{x},\mathbf{f})$, where $\mathbf{x}=(x_{1},x_{2},x_{3})$ and $\mathbf{f}=(f_{1},f_{2},f_{3})$ is defined as:
$\begin{cases}f_{1}=x_{1},\\ f_{2}=x_{1}\lor x_{3},\\ f_{3}=x_{2}\land x_{3}.\end{cases}$
Its state transition graph under the asynchronous update mode is shown in Figure .1. In this network, $x_{1}$ is an input node, meaning that its state remains constant during the evolution of the system. Let us consider a target configuration defined by $x_{2}=0$; that is, we want the network to settle into an attractor where $x_{2}$ set to 0, while the states of the other nodes are not constrained.
(0,0,0)(0,0,1)(0,1,0)(0,1,1)(1,0,0)(1,0,1)(1,1,0)(1,1,1)
1.Figure: State transition graph with highlighted attractor states of the Boolean network in under the asynchronous update mode.
If the input node $x_{1}=0$, then the target configuration is realised only by the attractor state $(0,0,0)$, which can be reached from the other attractor through a single perturbation:
$(0,1,1)\xrightarrow{i=3}(0,1,0)\xrightarrow{\text{evolution}}(0,0,0),$
where $\xrightarrow{i=3}$ denotes a change in state caused by flipping the value of node $x_{3}$, and $\xrightarrow{\text{evolution}}$ represents a sequence of transitions consistent with the natural dynamics of the network.
If $x_{1}=1$, there is only one attractor state that realises the target configuration, namely $(1,0,0)$. This state is reached through natural evolution from $(1,1,0)$, while a single perturbation of the input node is required to reach it from $(1,1,1)$:
$(1,1,1)\xrightarrow{i=1}(1,1,0)\xrightarrow{\text{evolution}}(1,0,0).$
-B Example: Attractor-target control with the switch of environmental conditions
We consider a BN with 5 nodes, i.e. $x_{1},x_{2},x_{3},x_{4},x_{5}$. The nodes $x_{1}$ and $x_{2}$ are designated as input nodes and the control objective is to achieve the target configuration $x_{5}=1$. Let us assume that the network is as follows.
  • For $x_{1}=0,x_{2}=*$, the network has two fixed-point attractors:
    $\{(0,0,0,0,0)\},\quad\{(0,0,1,1,0)\}.$
    In both cases $x_{5}=0$, so there is no attractor aligned with the target configuration.
  • For $x_{1}=1,x_{2}=1$, the attractors are:
    $\{(1,1,1,0,1)\},\quad\{(1,1,0,1,0)\}.$
  • For $x_{1}=1,x_{2}=0$, the attractor is:
    $\{(1,0,0,0,1)\}.$
Thus, there are three source attractors and two target attractors.
We consider $\{(0,0,0,0,0)\}$ to be the source attractor. Without perturbations of the input nodes $x_{1}$ and $x_{2}$, the system remains trapped in the configuration $x_{5}=0$. Therefore, perturbations of input nodes are necessary to change the environmental condition. There are two options: set $x_{1}=1,x_{2}=1$ or $x_{1}=1,x_{2}=0$. We consider both.
  • Suppose that the perturbation to state $(1,1,0,0,0)$ finally leads to the attractor:
    $\{(1,1,0,1,0)\}.$
    Since $x_{5}=0$, an additional intervention on $x_{3}$ is required to reach the target attractor, which is a CABEAN control path of length 1. Therefore, the total control strategy length in this case is 2.
  • The perturbation to state $(1,0,0,0,0)$ leads directly to the attractor:
    $\{(1,0,0,0,1)\},$
    which is a target attractor. Here, the total control strategy length is 1.
Consequently, the length of the reference control strategy obtained with CABEAN is 1 for this example.
-C Choosing iPASIP hyperparameters
The iPASIP hyperparameters should be set on a case-by-case basis in accordance with the available prior knowledge on the system under study, i.e. expected number and size of attractors. Additionally, for the setting of the $k_{1}$ hyperparameter (and $k_{2}$ by analogy) the information on the upper bounds on the size of the associated pseudo-attractors can be exploited as provided by the following theorem, which was originally proved in [ours_tcs].
Theorem .1.
For any attractor $\mathcal{A}$ of a BN under the asynchronous update mode, the size of the associated pseudo-attractor $\mathcal{PA}$ found by Step I-1 of the pseudo-attractor identification procedure with $k_{1}=k\%$ identification threshold is exactly upper bounded by
$|\mathcal{PA}|\leq\begin{cases}\frac{100}{k}-1,&100\text{ mod }k=0\text{ and }|\mathcal{A}|>\frac{100}{k}\\ \left\lfloor\frac{100}{k}\right\rfloor,&\text{otherwise}.\end{cases}$
Proof.
Since all states of $A$ are positive recurrent in the Markov chain underlying the BN restricted to $A$, it holds that $\mathbb{P}_{A}(s)>0$ for all $s\in A$. We start with the first case, where $100\text{ mod }k=0$ and $|A|>\frac{100}{k}$. A state $s$ is identified as a PA state by Step I if $\mathbb{P}_{A}(s)\geq\frac{k}{100}$. A pseudo-attractor associated with $A$ will be of maximum size whenever the stationary distribution maximises the number of states with $\frac{k}{100}$ probability. The maximum must be less than $\frac{100}{k}$. To show this, let $M$ denote the subset of states of $A$ with stationary distribution probability of $\frac{k}{100}$ and let assume that $|M|\geq\frac{100}{k}$. Then, $\sum_{s\in A}\mathbb{P}_{A}(s)=\sum_{s\in M}\mathbb{P}_{A}(s)+\sum_{s\in A\setminus M}\mathbb{P}_{A}(s)\geq\frac{k}{100}\cdot\frac{100}{k}+\sum_{s\in A\setminus M}\mathbb{P}_{A}(s)>1$, where the last inequality follows from the fact that all the states in $A\setminus M$ have non-zero stationary probabilities. Therefore, $\mathbb{P}_{A}$ cannot be a probability distribution. Hence, $|M|\leq\frac{100}{k}-1$. The upper bound is reached, for example, by any stationary distribution $\mathbb{P}_{A}$ such that the for the $\frac{100}{k}-1$ states in $M$ the probabilities are equal to $\frac{k}{100}$ and $\sum_{s\in A\setminus M}\mathbb{P}_{A}(s)=\frac{k}{100}$. This concludes the first case.
In the other case, let $\epsilon:=\frac{100}{k}-\left\lfloor\frac{100}{k}\right\rfloor$. If $|A|\leq\left\lfloor\frac{100}{k}\right\rfloor$, then the inequality holds since $PA\subseteq A$ by definition. The upper bound is reached if $|A|=\left\lfloor\frac{100}{k}\right\rfloor$ and, for example, the stationary distribution probabilities for |A|1|A|-1 states in $A$ are equal to $\frac{k}{100}$ and for the remaining state the probability is $\frac{k}{100}\cdot\left(1+\epsilon\right)$. If $|A|>\left\lfloor\frac{100}{k}\right\rfloor$ in the second case, then $\epsilon>0$, i.e. $\epsilon\in(0,1)$. The maximum possible number of states of $A$ with stationary probability equal $\frac{k}{100}$ must be less than $\left\lfloor\frac{100}{k}\right\rfloor+1$. To see this, let $M$ be the set of such states of $A$ and let assume that $|M|\geq\left\lfloor\frac{100}{k}\right\rfloor+1$. Then, $\sum_{s\in A}\mathbb{P}_{A}(s)>\frac{k}{100}\cdot\left(\left\lfloor\frac{100}{k}\right\rfloor+1\right)>\frac{k}{100}\cdot\left(\left\lfloor\frac{100}{k}\right\rfloor+\epsilon\right)=1$. This contradicts the fact that $\mathbb{P}_{A}$ is a probability distribution on $A$. Finally, we proceed to show that the upper bound can be reached also in this scenario. This is the case, for example, for a stationary probability distribution $\mathbb{P}_{A}$ on $A$ where $\left\lfloor\frac{100}{k}\right\rfloor$ states of set $M$ have probabilities $\frac{k}{100}$ and all the states in $A\setminus M$ have non-zero probabilities such that $\sum_{s\in A\setminus M}\mathbb{P}_{A}(s)=\epsilon$. This is possible since $\epsilon>0$. ∎